When the price of Dragon fruit was Rs. 100 per kg, 200kgs of Dragon fruit were demanded in Bandra market per day. Now the Dragon fruit price has increased to Rs. 110 per kg as a result of which the quantity demanded reduced to 160 kg per day. Calculate the price elasticity of demand by using the mid - point method


1
Expert's answer
2021-05-24T13:57:51-0400

price elasticity of demand"=\frac {\frac{Q_2-Q_1}{(Q_2+Q_)/2}}{\frac{P_2-P_1}{(P_2+P_1)/2}}"


"=\frac {\frac{160-200}{(160+200)/2}}{\frac{110-100}{(110+100)/2}}"


"=\frac{-40}{180}\div\frac{10}{105}"


"=\frac{-40}{180}\times \frac{105}{10}=-2.33"

"=-2.33"